The Malayo Colleges Mindanao recently conducted a study analyzing the privacy management habits of teen internet users. In a group of 118 teens, 108 reported having more than 100 followers on Twitter. Find the upper bound of a 90% confidence interval for the true proportion of teens who would report having more than 100 followers on Twitter. 



1
Expert's answer
2022-03-08T12:34:06-0500

"p=\frac{108}{118}=0.9153."

Upper bound "=0.9153+1.645\sqrt{\frac{0.9153(1-0.9153}{118}}=0.9575."